OpenStudy (anonymous):

|| I WILL GIVE MEDAL AND FAN || Ashley solved the exponential equation 3x+1 = 15 and her work is shown below. What is the first step she did incorrectly?

OpenStudy (anonymous):

Step 1: log 3^x+1= log15

OpenStudy (loveyourself99):

log?

OpenStudy (anonymous):

Step 2: (x + 1)log 3 = log15 Step 3: x + 1 = log 15 over log 3 Step 4: x+1 = 2.708050 over 1.098612 Step 5: x + 1 = 1.418858 Step 6: x = 0.418858

OpenStudy (anonymous):

I'm not totally postive what a log is for Algebra 2

OpenStudy (loveyourself99):

you take 15 -1 that gives you 14. you take 14 divided by 3

OpenStudy (loveyourself99):

that gives you 4.66666666667

OpenStudy (anonymous):

So Step 4 is wrong @loveyourself99

OpenStudy (loveyourself99):

yes very wrong

OpenStudy (anonymous):

Thank you for your help and I kinda understand it better now @loveyourself99

OpenStudy (loveyourself99):

you're welcome i love this kind of math